OpenStudy (anonymous):

AD=3 centimeters, so AB= ___ centimeters.

OpenStudy (anonymous):

 @souvik last question can you help me

OpenStudy (souvik):

ABcos60=AD find AD

OpenStudy (anonymous):

but we already have ad its 3cm

OpenStudy (souvik):

sorry...find AB

OpenStudy (anonymous):

how do i do that im so confused :/

OpenStudy (souvik):

AB=AD/cos60

OpenStudy (anonymous):

i got 0.989 is that right

OpenStudy (souvik):

cos60=0.5 AB=6 cm

OpenStudy (anonymous):

how did you do that i put it correct in the calculator and i gave me 1 then i did it again and i got 0.989

OpenStudy (souvik):

cos60=?

OpenStudy (anonymous):

yahh i would put that then i would put 60 put it didnt giveme anything

OpenStudy (souvik):

first put 60 then cos...hope this time works

OpenStudy (anonymous):

okay i did that then what do i do

OpenStudy (souvik):

so cos60=0.5 now AB=AD/0.5=3/0.5=6 cm

OpenStudy (anonymous):

okay yes now i got it i thought you where suppost to put cos first then the 60 thank you

OpenStudy (souvik):

here in India we put cos first and then 60....:)

OpenStudy (anonymous):

ohhh okay isnt it already late for you over there?

OpenStudy (souvik):

12:30 AM here..

OpenStudy (anonymous):

really wow over here in italy its 9:00 pm well i guess ill let you go soyou can sleep just one moer question how do ou say bye in india?

OpenStudy (souvik):

there are many language in India...in Bengali..we use bye-English word.. but informally "ta-ta" , formally "biday" .... good night="suvo rattri"

OpenStudy (souvik):

in Hindi...bye='albida'